Grade 11Mechanics

the moment of inertia of a fly wheelmaking 300 revolutions per minute is 0.3kgm2.find torque required to bring it to rest in 20 sec?

zeeshan siddiqui
10 Years agoGrade 11

1 Answer

arun
ApprovedApproved Tutor Answer10 Years ago
it is given that the angular velocity of fly wheel is 300 rpm = 10\pi rad/s
and t= 20 sec
let initial angular velocity be \omega and final be \varpi
so, \omega = 10\pi , \varpi = 0
let angular acceleration is -\alpha
using first equation  of rotational motion,
\varpi = \omega - \alpha t
10\pi = 20\alpha
\alpha=\frac{\pi}{2}
and \tau =I\alpha \Rightarrow \tau = 0.3\times \frac{\pi}{2}
\tau =0.15\times 3.14 \Rightarrow \tau= 0.47 Nm